4. Find the missing measure of each triangle.
A= 50, B= 70 c= ?

Answer by checkley77(12844)

IF c= hypotenuse & this is a right triangle then:
50^2+60^2=c^2
2500+3600=c^2c^2
6,100=c^2
c=sqrt6,100
c=78.102496 ans.